3(8 points) Consider the expansion of 1.00 mole of (ideal) Ne from 2.00 atm at 75.00 C; the volume is doubled in the process. Find q, w, AH and the final pressure and temperature for a) reversible adiabatic expansion b) reversible isothermal expansion c) reversible constant pressure expansion d) irreversible adiabat against 0.500 atm external pressure
Answer #1

a) for reversile adiabatic expansion,q=0 , W= nR*(T2-T1)/ (Y-1), Y= CP/CV= ratio of specific heats of Ne=1.67

n= no of moles of gas

R= 8.314 Joules/mole.K, Temperature( T2) at the end of expansion process is

T2/T1= (V1/V2) R/CV= (V1/V2) (Y-1)

T1= 75+273= 348 K, V1= V and V2=2V

T2= 348*(1/2)(1.67-1)=218.7 K

hecne work done = 1*8.314*(218.7-348)/0.67=-1604 Joules, work is -ve since system does work on the surroundings

since from 1st law of thermodynamics, deltaU = q+W

deltaU= change in internal energy= -1604 joules for neonm CP= 20.78 J/mole.K

deltaH= n*CP*temperature difference= 1*20.78* (218.7-348)= -2687 Joules

since the expansion is adiabatic, P1V1Y= P2V2Y

2*VY= P2*(2V)Y

P2 = 2*(1/2)Y = 0.63 atm

b) for isothermal compression of an ideal gas, deltaU=0, deltaH=0 and deltaU=W= nRTln(V1/V2)

=1*8.314*348*ln(1/2)= -2005 joules

c) for constant pressure ( isobaric expansion)

from gas law P1V1/T1= P2V2/T2

gvien V2= 2V1

P1= 2atm, T1= 348K, P2= 2atm, V1=V and V2=2V

T2= (V2/V1)*T1=2*348= 696K

work done = -nR*(T2-T1)=-8.314*(696-348)= -2893 joules

deltaH= nCP*dT=1*20.78*(696-348)= 7231 Joules

deltaU= nCv*(T2-T1)= (20.78-8.314)*(696-348)=4338 joules

q=deltaH=7231 joules

4. work done= -Pext.dV, dV= volume change, volume initially, V1= nRT/P= 1*0.0821 L.atm/mole.K* 348/2= 14.28 L

volume after expansion =2*14.28= 28.56L

work done= -0.5*(28.56-14.28)L.atm. 101.3J/L.atm =-723 joules

for adiabatic, q=0 and deltaU= -723 joules